Evaluate

log2 16 and log7 1/49

Answer by edjones(7311)

y=log[a]X if X=a^y
y=log[2]16 if 2^y=16
2^4=16
log[2]16=4
.
y=log[7]1/49 if 7^y=1/49
7^-2=1/49
log[7]1/49=-2
.
Ed